OpenStudy (flygirlmandi):

find the missing length for a right triangle with hypotenuse of 18 and leg length of 6. round to the nearest whole number if necessary. A-17 B-324 C-19 D-36 Thanks :)

OpenStudy (mysticfury):

formula is. a2 + b2 = c2 the hypotenuse is always C and the leg is always A or B dosent matter which one plug in the numbers... 6^2 + B^2 = 18^2 simplify 36 + b^2 = 324 subtract 36 from both sides 36-36 is 0 ... 324 - 36 is 288 b x b is b so we got b = 288 now find the square root of 288 16.97 rounded its 17 so your answer is 17 :D

OpenStudy (flygirlmandi):

thanks for explaining it, im really bad at math..lol XD

OpenStudy (mysticfury):

its okay, i didnt get this theory either, but i just did it a little while ago :D happy to help out.. you understand it a little better now? its real easy once you get the hang of it... and believe me if i say somethings easy its easy, im the kind of person thats sooo bad at math , some stuff people say is so easy is sooooooo difficult to me, even when i know how to do it XD

OpenStudy (flygirlmandi):

:D

OpenStudy (mysticfury):

if you need anymore help just tag me :)